OpenStudy (anonymous):

A company has determined that 1% of its widgets are defective. If a customer received 2 widgets, what is the probability that they are both defective? 1/10 1/10,000 1/100 1/1,000

OpenStudy (anonymous):

Assume independence because there are probably a lot of widgets. 1/100 x 1/100 = 1/10,000

OpenStudy (anonymous):

How did you get the 1/100?

OpenStudy (anonymous):

From 1% 1% = 0.01 = 1/100

OpenStudy (anonymous):

Oh, okay. Thank you for helping me :)

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @k3nzi14 and you're welcome!